Notes from AW; apologies for misinterpretations of what was said.

Brian Kent, Emerson, Nyman, Hills, Wiklind, Peck, Wootten, Sawada, Paola, eelco, Klaassen, Hogerheijde,

OSF interferometer is up. Weather is combatative. Monthly report--one which goes to the Board, with warning. 8ns delay problem. Often a long delay in a problem getting fixed. Problem with DV01 set on its 3point mounting; friction caused a deformation in the base solved by picking it up and setting it down to release the friction. Many of the issues are software. Correlator often times out when writing to archive. Problem reproduced in simulation in Cville but we await a fix. Anxious to move antennas to shorter baselines. That was to occur 1st of January, now estimate 26-27 Feb. Getting the new power for the first time, with safety procedures and staffing them has been a challenge. Two antennas should be on the ACA pads by the end of the month. We will then run a short time with these two and the one left ~1km away. We hope to test phase correction on the long baselines.

Three new people arrived in January, a few more on Ops side, one more next week. We are pretty well up to planned staffing. ARC support scientists have come down also.

There is a wiki page with more detailed tasks. On the groups some external people have been involved in these tasks. If those on line wish to participate please get in touch.

Cost saving exercise. Both construction and operations costs have posed problems. Items spoken about last time including reducing subarrays and doing line length correction in software have been dropped as savings were small and there was risk. There are no other cost saving issues we need to discuss. there will be a list before the Board meeting. Paola: Ops? Nyman: We are updating the Ops plan. The budget has been pressured by needs that have evolved so we were asked to come up with a modified budget which could bring the 2015 onwards budget down to the level of the approved version. Power is now $9.5M in gas turbines compared to grid estImate of $3.5M. It is difficult to find such savings. Another iteration is due today. Deadline is end of this month to submit to ALMA budget committee. No Bolivian winter shutdown, cycle skipping some configurations is being considered. Eight going out and eight others going in with a one year cycle, saves two transporter crews. DSO composition change also considered, 13 AOds and 9 fellows, changed fellows to data analysts. Originally 4 8 hour shifts per day to allow ACA parallel ops, by 2015 we don't expect that so reduce one shift. thus need only 9 astronomers and 6 data analysts. Paola--thanks for mentioning this; reducing astronomers and trying to eliminate fellows is a real problem. Nyman: work is ongoing. Do we really need the 7 MW. Paola: I find this scary. REH: Scientists are used to fine tuning; others are not. There is an issue of balance here. I hope that those involved are arguing these points as strongly as possible. Nyman: There will be a result with a complete picture by the end of the month, with all departments including facilities. REH: We will need to look at the total picture. Wright: If we want people involved, they need the data.

Change requests. Relaxation of band 6 polarization. Has been in works for a long time. Higher levels of cross pol measured. Seems to be a sum of contributions of lots of small parts to the whole. REH feels the estimate is pessimistic. It is thought that the requirement might be met in lower part of band. Band 6 frequency coverage. Mixers and IF amps will provide reasonable performance with 5 GHz IF. Excess noise added by WCA in LO system is not solved. A revised version of the multiplier chain might be needed. The WCAs for other receivers do not have this problem.

FE CDR. Linearity and how nonlinearity has been discussed, along with differential phase stability. Our problem at the moment is having front ends, the one to be solved at the moment.

Calibration device. JGM. One issue was measuring saturation. REH: We are investigating the ACDs but they don't move reliably under software control and has been a subject of much discussion. Has the oxygen sounder order gone in.

WVR report.

IR Camera We should get on with this.

Splatalogue. Brian--test server at NCSA. Small class library in php. An idl example will allow a return of an xml table over a frequency range. NCSA will package in an installer which will allow it to be used at other sites including JPL.

Upcoming meetings.

Reasonable to ask--lots going on down here on commissioning side. Purpose of this telecon is to keep people who are interested in touch with what is going on. Pretty small attendance. Wright: thinking about it. This is a narrative meeting, for specific items it is best to have specific meetings. REH: Some items, like the operations let us know. Change requests are now less important.

-- AlWootten - 2010-03-16
Topic revision: r1 - 2010-03-16, AlWootten
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback